OpenStudy (anonymous):

2l ogx + logy - 3 logz

OpenStudy (anonymous):

it would be log(x^2y/2z^3)

OpenStudy (anonymous):

Accdg to the laws of logarithm 2logx = logx^2 3logz = logz^3 sum of 2 log is the log of its product: log(x^2 y)- logz^3 difference of 2 log is the log of a quotient therefore the simplified form is log(x^2y)/logz^3

OpenStudy (anonymous):

can you answer this 9^(x-1)=27

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

We can use the logarithmic equations (x-1) log 9 = log 27 x-1 = log 27/log9

OpenStudy (anonymous):

you can continue from there

OpenStudy (anonymous):

What would my answer be?

OpenStudy (anonymous):

Oops the answer to the first question is log{(x^2y)/z^3}

OpenStudy (anonymous):

are you allowed to use calculator to find the value of log 9 and log 27?

OpenStudy (anonymous):

i am not sure, thanks for the help